OpenStudy (anonymous):

A G.P consist of 2n terms. If the sum of the terms occupying the odd places is S(1) and that of the terms occupying the even places is S(2) , then find common ratio of the progression

OpenStudy (anonymous):

I am stuck

OpenStudy (anonymous):

We can see here the odd places and even places then apply the sum of infinite terms in a G.P formula after that what should i do

hartnn (hartnn):

what is S(1) ?

OpenStudy (anonymous):

Oh there is no value given

hartnn (hartnn):

i mean, what does it represent ? or is it just some random constant like m,n ...

OpenStudy (anonymous):

It represents sum of terms occupying the odd places. Yeah constant

OpenStudy (anonymous):

s(1) is coming a/1-r^3

hartnn (hartnn):

they want common ratio in terms of S1 and S2 , right ?? a, ar, ar^2, ar^3 ,... series odd : a, ar^2,ar^4 ... common ratio = r^2 S1 = a/(1-r^2 ) series even, ar, ar^3 , ar^5.... S2 = ar /(1-r^2)

hartnn (hartnn):

this turned out to be simple

hartnn (hartnn):

just divide those

hartnn (hartnn):

nopes, r = S2/S1

hartnn (hartnn):

S2 = ar /(1-r^2) S1 = a/(1-r^2 ) a and 1-r^2 gets cancelled

OpenStudy (anonymous):

yes s2/s1

OpenStudy (anonymous):

sorry

hartnn (hartnn):

np :)